glizxzy:

Find the derivative of y=√x+83√x-24√x .

glizxzy:

@Aqual

glizxzy:

@yeyedontplay

glizxzy:

@dericawarren

Aqual:

isnt the derivative like a slope?

glizxzy:

Derivatives are a fundamental tool of calculus. ... The derivative of a function of a single variable at a chosen input value, when it exists, is the slope of the tangent line to the graph of the function at that point. The tangent line is the best linear approximation of the function near that input value. is the exact defenition from my notes

Noluvpablo:

are we solving for y?

glizxzy:

Remember that you’ll need to convert the roots to fractional exponents before you start taking the derivative. Here is the rewritten function. y=x1/2+8x1/3−2x1/4 The derivative is,dy/dx=1/2x−1/2+8/3x−2/3−1/2x−3/4